OpenStudy (anonymous):

Help i always give metals ;) When simplified, what is the value of square root of negative 128 end square root, if i^2 = −1?

OpenStudy (campbell_st):

well making the substitution you get -128 = 128i^2 which can be written as 2 x 64 x i^2 now you should be able to simplify it

OpenStudy (anonymous):

hmm is it 

Nnesha (nnesha):

\[\sqrt{-128}\] whenever u see negative sign separate that from the number\[\sqrt{-1 \times 128}\] now use the fact \[i=\sqrt{-1}\] and as mentioned above 128 can be written as 2 times 64\[\sqrt{-1 *2*64} \rightarrow \sqrt{-1} *\sqrt{64}*\sqrt{2}\]